OpenStudy (anonymous):

If f(x) = exg(x), where g(0) = 6 and g'(0) = 4, find f '(0).

OpenStudy (anonymous):

exg?

OpenStudy (anonymous):

\[ f(x) = e^xg(x) \]This?

OpenStudy (anonymous):

Maybe \[ f(x) = e^{xg(x)} \]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Okay, then \[ f'(x) = e^xg(x) + e^xg'(x) \]

OpenStudy (anonymous):

\[ f'(0) = e^0g(0)+e^0g'(0) \]

OpenStudy (anonymous):

Can you finish it off?

OpenStudy (anonymous):

yes, thank you